# Adaptive Accelerated Composite Minimization

### v1

@@ Abstract.
The choice of the stepsize in first-order convex optimization is typically based on the smoothness constant and plays a crucial role in the performance of algorithms. Recently, there has been a resurgent interest in introducing adaptive stepsizes that do not explicitly depend on smooth constant. In this paper, we propose a novel adaptive stepsize rule based on function evaluations (i.e., zero-order information) that enjoys provable convergence guarantees for both accelerated and non-accelerated gradient descent. We further discuss the similarities and differences between the proposed stepsize regimes and the existing stepsize rules (including Polyak and Armijo). Numerically, we benchmark the performance of our proposed algorithms with the state-of-the-art literature in three different classes of smooth minimization (logistic regression, quadratic programming, log-sum-exponential, and approximate semidefinite programming), composite minimization ($\ell_{1}$ constrained and regularized problems), and non-convex minimization (cubic problem).
Keywords: Adaptive stepsize, first-order methods, composite convex optimization
@@ 1. Introduction
Thanks to its simple implementation and applicability, gradient descent (GD) is perhaps the most popular algorithm in convex optimization [5, 14]. The main challenge in using gradient descent is the choice of the right stepsize, which has a considerable impact on the convergence speed. Several works have studied this choice, yet either the fastest possible convergence rate cannot be guaranteed or an expensive linesearch at each iteration is needed [37, 24, 34, 15]. In this paper, we address these issues by introducing a novel and simple linesearch method.
@@
(i) GD with constant stepsize [34]: The simplest stepsize rule is the constant $\lambda_{k}=1/{2L}$, where $L$ is the so-called smoothness parameter (Lipschitz continuity of $\nabla f(x)$. The convergence rate of the suboptimality $f(x_{k})-f(x^{*})$ with the constant stepsize is $\mathcal{O}(k^{-1})$.
(ii) GD with exact linesearch [34]: Another classical stepsize is to find the optimal stepsize minimizing the function $\phi_{k}$ along the direction $d_{k}$ is $\lambda_{k}=\arg\min_{\lambda}\phi_{k}(\lambda)$ which involves a scalar optimization problem that requires evaluation of the original function $f$.
(iii) GD with approximated stepsize [34]: Due to the complexity of finding the stepsize by solving the exact linesearch minimization, alternative stepsize rules have been proposed that only require ensuring an inequality such as $\phi_{k}(\lambda)\leq\phi_{k}(0)+c\phi_{k}^{\prime}(0)$ for some predefined constant $c>0$. Among others, Armijo [2], Wolf, and strong Wolf linesearch [34] are backtracking linesearch methods falling into this category.
@@
The suboptimality convergence rate of these algorithms is $\mathcal{O}(k^{-1})$. There are also several stepsize rules approximating the smooth constant locally based on the zero-order oracle’s information (i.e., the objective function evaluation) [36, 24, 3, 4]. In the following, we review two important stepsize rules that fall into this category:
(iv) GD with the Polyak stepsize [36]: One of the classic adaptive stepsize rules is the Polyak stepsize, which uses the zero-order oracle’s information at each iteration to locally approximate the smooth constant of the objective function:
@@
Although the Polyak stepsize offers optimal convergence rates for GD (1) when the update direction is set to $d_{k}=-\nabla f(x_{k})$ [16], it cannot be applicable in problems where the optimal function value $f(x^{*})$ is not available. We will get back to the Polyak stepsize and its relation to our proposed rule in Subsection 2.3 (see equations (18)).
(v) GD with adaptive stepsize [24]: Recently, the next adaptive stepsize has been proposed based on the idea of approximating the smoothness constant locally:
@@
The second term on the left-hand side of (5) represents this local approximation as the inverse of the smoothness constant, while the first term ensures a required convergence rate of this parameter. The theoretical sub-optimality convergence rate for the stepsize rule (5) is $\mathcal{O}(k^{-1})$, however, the adaptive (i.e., state-dependent) nature of (5) has the advantage of (a) lifting the knowledge of the smoothness constant for the objective function, and (b) improving the practical convergence rate when the local smoothness is smaller than the global constant one. Next, we consider methods with a dynamic direction $d_{k}$ in (1). The update rules in this class of methods have an extra term, which is called momentum, and allows the descent direction to have some inertia in the search space [37].
(vi) AGD with constant stepsize [29]:
@@
(vii) AGD with adaptive stepsize [24]: A heuristic version of adaptive stepsize (5) for accelerated methods is also proposed in [24, Section 3]
@@
where $f\colon\mathcal{X}\to\mathbb{R}$ is a convex and smooth function, and $h\colon\mathcal{X}\to\mathbb{R}$ is convex and possibly non-smooth but prox-friendly (technical details are postponed to Section 2). There are several methods for convex composite minimization in the existing literature. Among them, subgradient descent and mirror descent are two well-known classic methods that suffer from slow convergence rate $\mathcal{O}(k^{-1/2})$ [28, 7, 42, 40]. Following the seminal works by Nesterov, one can exploit the structure of the nonsmooth part in the objective function and deploy a first-order accelerated method. ISTA and FISTA [8], which are extensions of GD and NAGD, respectively, offer a faster convergence rate compared to subgradient and mirror descent. However, these methods require knowledge of the smooth constant associated with the smooth part, $f$ in (6). Additionally, there are various backtracking techniques available to approximate the smooth constant, but they often tend to be conservative, resulting in larger values than the actual smooth constant. This conservatism can impact the convergence speed. Recently, the authors in [39, 23] propose an algorithm that enjoys the local smoothness of $f$ to determine the stepsize. However, despite its simplicity and efficacy, the convergence guarantee is still the suboptimal rate of $\mathcal{O}(k^{-1})$. Inspired by this, we develop an algorithm for convex composite minimization problems (6) with the best theoretical convergence rate $\mathcal{O}(k^{-2})$ while maintaining a reasonable level of computational complexity.
@@
Contributions. Following the same spirit of the adaptive stepsize rules, this study proposes a novel stepsize rule without the knowledge of the global smoothness constant for solving convex composite minimization problems. Using a Lyapunov-based argument, we show that the proposed rule enjoys the optimal worst-case complexity bound of $\mathcal{O}(k^{-1})$ and $\mathcal{O}(k^{-2})$ in both cases of nonaccelerated and accelerated algorithms, respectively. These results are developed for a general class of composite convex functions and are optimal in the sense that they match the theoretical lower bounds of the class of first-order algorithms for this class of functions. Our convergence results are summarized in Table 1 where our update direction $d_{k}=-G^{f}_{\lambda h}(x_{k})$ is our gradient mapping (see the definition (10) for more details concerning the gradient mapping). Let us note that the results in Table 1 are reported for the general composite case (6). If the nonsmooth term is $h(x)=0$, then the gradient mapping reduces to $G^{f}_{\lambda h}(x)=\nabla f(x)$. In the rest of the introduction, we briefly discuss our proposed stepsize rule in this special case of a smooth convex function and provide a geometrical illustration in comparison with the existing rules reviewed above.
@@
where the function $\phi_{k}$ is defined in (2) and visualized in Figure 1a. Recall that in the case of non-accelerated methods, the update direction is simply $d_{k}=-\nabla f(x)$. We emphasize that the proposed linesearch above only requires additional zeroth-order information, namely the objective function evaluation, while the first-order information is the gradient of the function which is already computed in the previous step. Figure 1b illustrates a geometric interpretation of the proposed stepsize rule (7), together with the existing literature reviewed in the preceding section. For further geometrical interpretation, particularly in case of accelerated methods, we refer to [13].
@@
[…]
@@
Furthermore, $f$ is locally smooth if it is smooth over any compact set of its domain, see [32, Chap. 2] for more details. The prox operator of a convex function $h\colon\mathcal{X}\to\mathbb{R}$ is defined as
@@ 2.1. Preliminaries
We first proceed with some assumptions and lemmas that will be used throughout the paper. We note that the classical Cauchy-Schwartz and convexity inequalities are the two useful tools in our analysis and most of the results build on the seminal works by Nesterov [29] and Polyak [37]. The following assumptions holds throughout this study.
@@
$G^{f}_{\lambda h}(\hat{x})=0$ $\iff\hat{x}=\text{prox}_{\lambda h}\left(\hat{x}-\lambda\nabla f(\hat{x})\right)\iff(\hat{x}-\lambda\nabla f(\hat{x}))-\hat{x}\in\lambda\partial h(\hat{x})$
@@
$r=\arg\min_{r}h(r)+\dfrac{1}{2\lambda}||r-t||^{2}\iff 0\in\partial{h(r)}+\dfrac{1}{\lambda}(r-t)\iff t-r\in\lambda\partial{h(r)}.$ (12)
@@ 2.2. Non-accelerated composite minimization
$\begin{cases}\lambda_{k}=\max\limits\left\{\lambda\in\mathbb{R}_{+}\,|\,\,\phi_{k}(2\lambda)\leq\phi_{k}(\lambda)-\lambda\langle G^{f}_{\lambda h}(x_{k}),\nabla f(x_{k})\rangle+\dfrac{\lambda}{2}\|G^{f}_{\lambda h}(x_{k}))\|^{2}\right\},\\ d_{k}=-G^{f}_{\lambda_{k}h}(x_{k}).\\ \end{cases}$ (Alg. 1)
Then, we have the uniform stepsize lower bound $\lambda_{k}\geq{1}/{2L}$ where $L$ is the smoothness constant of $f$. Moreover, we have the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D}{k}$ where $D$ is a constant that only depends on the initial condition $x_{0}$, the optimal solution $x^{*}$, and the smoothness constant $L$.
@@
First, to show that $\lambda_{k}$ is bounded from below by ${1}/{2L}$, we just need to show that the condition in the stepsize rule always is satisfied by $\lambda_{k}=1/{2L}$. Considering the smoothness of $f$, we can write the inequality (8b) for the pair $(x,x^{+})$ where $x^{+}=x-\lambda G^{f}_{\lambda h}(x)$, which arrives at the two inequalities
@@
which is the same as inequality condition in the stepsize rule when $\lambda=1/{2L}$, demonstrating that the stepsize rule condition in Alg. 1 is always satisfied by $\lambda=1/{2L}$. Next, we prove the convergence of Alg. 1. By putting $z=x_{k}$ in the inequality (iii) we have
$F(x_{k+1})-F(x_{k})\leq-\dfrac{\lambda_{k}}{2}\|G^{f}_{\lambda h}(x_{k})\|^{2}$ (15)
Telescoping the above inequalities from $k=0$ to $k=\infty$ and using the fact that $\lambda_{k}\geq 1/2L$ gives us
$F(x^{\infty})-F(x_{0})\leq-\dfrac{1}{4L}\sum_{i=0}^{\infty}\|G^{f}_{\lambda h}(x_{i})\|^{2}$
@@
The above approximate solution is guaranteed to satisfy $\lambda_{k}\geq\dfrac{C}{2L}$.
@@ 2.3. Non-accelerated smooth minimization
Let the function $F$ in (6) be a smooth (i.e., $h(x)=0$). Then, the stepsize rule in Alg. 1 reduces to (7) while ensuring Theorem 2.3 results of the uniform stepsize lower bound $\lambda_{k}\geq{1}/{2L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D}{k}$.
@@
We close this subsection by highlighting a common feature shared between the Polyak stepsize (4), Armijo stepsize rule (3a), and our proposed stepsize rule (7). These three stepsizes, denoted by $\lambda$, fall into the interval
$\dfrac{1}{2L}\leq\lambda\leq c\,\dfrac{f(y_{1})-f(y_{2})}{\|\nabla f(x_{k})\|^{2}},$ (18a)
@@
The common lower bound $1/2L$ in (18a) is well known for Polyak and Armijo, and it is also shown for our proposed rule in Theorem 2.3. Notice that the optimal choice of the stepsizes is the upper bound of the admissible interval (18a). It is also important to note that in the case of Armijo (3a) and our proposed choice (7), this upper bound also depends on $\lambda$ (see (18b)), making the stepsize rule implicit. Finally, we wish to draw attention to the choice of the upper bound between Armijo and our proposed approach in (18a): One can inspect that the proposed rule (7) looks one step ahead of Armijo in the direction of $\nabla f$. This may explain why the proposed rule (7) performs numerically better, as it can approximate the curvature of the function $f$ by “looking one step further” than Armijo. Finally, it is also worth noting that the recent work [24] offers an adaptive stepsize rule with the same lower bound $1/2L$ and an explicit upper bound (i.e., independent of $\lambda$) that depends on the gradient of two successive iterations.

### v2

# A New Lineserach for Accelerated Composite Minimization

@@ Abstract.
The choice of the stepsize in first-order convex optimization is typically based on the smoothness constant and plays a crucial role in the performance of algorithms. Recently, there has been a resurgent interest in introducing adaptive stepsizes that do not explicitly depend on smooth constant. In this paper, we propose a novel linesearch stepsize rule based on function evaluations (i.e., zero-order information) that enjoys provable convergence guarantees for both accelerated and non-accelerated gradient descent. We further discuss the similarities and differences between the proposed stepsize regimes and the existing stepsize rules (including Polyak and Armijo). We numerically benchmark the performance of our proposed algorithms against state-of-the-art methods across three major problems classes of (1) smooth minimization (logistic regression, quadratic programs, log-sum-exponential, and smooth max-cut relaxation) (2) composite minimization ($\ell_{1}$-regularized least-squares, $\ell_{1}$-constrained least-squares, and $\ell_{1}$-regularized logistic regression), and (3) non-convex minimization (cubic minimization). These classes include a wide range of operations research and management applications such as portfolio optimization, discrete choice models, sparse classification and feature selections, high-order optimization and trust-region subproblems.
Keywords: Adaptive stepsize, linesearch rules, first-order methods, composite convex optimization
@@ 1. Introduction
Thanks to its simple implementation, scalability and applicability, gradient descent (GD) is arguably the most popular algorithm to efficiently solve large-scale optimization problems, which finds a wide range of applications in operations research [im2025stochastic, xu2021first, ho2018online], machine learning [Battiti1992, lan2020first], control and system identification [hardt2018gradient, fazel2013hankel]. The main challenge in using gradient descent is the choice of the right stepsize, which has a considerable impact on the convergence speed. Several works have studied this choice, yet either the fastest possible convergence rate cannot be guaranteed or an expensive linesearch at each iteration is needed [Polyak1969, Malitsky2020, Nocedal2006, Drori2014]. In this paper, we address these issues by introducing a novel and simple linesearch method.
@@
(i) GD with constant stepsize [Nocedal2006]: The simplest stepsize rule is the constant $\lambda_{k}=1/{L}$, where $L$ is the so-called smoothness parameter (Lipschitz continuity of $\nabla f(x)$. The convergence rate of the suboptimality $f(x_{k})-f(x^{*})$ with the constant stepsize is $\mathcal{O}(k^{-1})$.
(ii) GD with exact linesearch [Nocedal2006]: Another classical stepsize is to find the optimal stepsize minimizing the function $\phi_{k}$ along the direction $d_{k}$ is $\lambda_{k}=\arg\min_{\lambda}\phi_{k}(\lambda)$ which involves a scalar optimization problem that requires evaluation of the original function $f$.
(iii) GD with approximated stepsize [Nocedal2006]: Due to the complexity of finding the stepsize by solving the exact linesearch minimization, alternative stepsize rules have been proposed that only require ensuring an inequality such as $\phi_{k}(\lambda)\leq\phi_{k}(0)+c\phi_{k}^{\prime}(0)$ for some predefined constant $c>0$. Among others, Armijo [Armijo1966], Wolf, and strong Wolf linesearch [Nocedal2006] are backtracking linesearch methods falling into this category.
@@
The suboptimality convergence rate of these algorithms is $\mathcal{O}(k^{-1})$. There are also several stepsize rules approximating the smooth constant locally based on the zero-order oracle’s information (i.e., the objective function evaluation) [poliak1987introduction, Malitsky2020, barre2019polyak, barre2020complexity]. In the following, we review two important stepsize rules that fall into this category:
(iv) GD with the Polyak stepsize [poliak1987introduction]: One of the classic adaptive stepsize rules is the Polyak stepsize, which uses the zero-order oracle’s information at each iteration to locally approximate the smooth constant of the objective function:
@@
Although the Polyak stepsize offers optimal convergence rates for GD (1) when the update direction is set to $d_{k}=-\nabla f(x_{k})$ [hazan2019revisiting], it cannot be applicable in problems where the optimal function value $f(x^{*})$ is not available. We will get back to the Polyak stepsize and its relation to our proposed rule in Subsection 2.3 (see equations (18)).
(v) GD with adaptive stepsize [Malitsky2020]: Recently, the next adaptive stepsize has been proposed based on the idea of approximating the smoothness constant locally:
@@
The second term on the left-hand side of (5) represents this local approximation as the inverse of the smoothness constant, while the first term ensures a required convergence rate of this parameter. The theoretical sub-optimality convergence rate for the stepsize rule (5) is $\mathcal{O}(k^{-1})$, however, the adaptive (i.e., state-dependent) nature of (5) has the advantage of (a) lifting the knowledge of the smoothness constant for the objective function, and (b) improving the practical convergence rate when the local smoothness is smaller than the global constant one. Next, we consider methods with a dynamic direction $d_{k}$ in (1). The update rules in this class of methods have an extra term, which is called momentum, and allows the descent direction to have some inertia in the search space [Polyak1969].
(vi) AGD with constant stepsize [Nesterov1983]:
@@
(vii) AGD with adaptive stepsize [Malitsky2020]: A heuristic version of adaptive stepsize (5) for accelerated methods is also proposed in [Malitsky2020, Section 3]
@@
where $f\colon\mathcal{X}\to\mathbb{R}$ is a convex and smooth function, and $h\colon\mathcal{X}\to\mathbb{R}$ is convex and possibly non-smooth but prox-friendly (technical details are postponed to Section 2). There are several methods for convex composite minimization in the existing literature. Among them, subgradient descent and mirror descent are two well-known classic methods that suffer from slow convergence rate $\mathcal{O}(k^{-1/2})$ [nemirovskij1983problem, Beck2003, yurtsever2018conditional, yang2018rsg]. Following the seminal works by Nesterov, one can exploit the structure of the nonsmooth part in the objective function and deploy a first-order accelerated method. ISTA and FISTA [Beck2009], which are extensions of GD and NAGD, respectively, offer a faster convergence rate compared to subgradient and mirror descent. However, these methods require knowledge of the smooth constant associated with the smooth part, $f$ in (6). Additionally, there are various backtracking techniques available to approximate the smooth constant, but they often tend to be conservative, resulting in larger values than the actual smooth constant. This conservatism can impact the convergence speed. The authors in [vladarean2021first, malitsky2020golden] propose an algorithm that enjoys the local smoothness of $f$ to determine the stepsize. However, despite its simplicity and efficacy, the convergence guarantee is still the suboptimal rate of $\mathcal{O}(k^{-1})$. More recently, the authors in [li2023simple] propose a fully adaptive stepsize with momentum acceleration and a convergence rate of $\mathcal{O}(k^{-2})$. However, the algorithm requires several hyperparameters, which may affect the performance in practice. Inspired by this, we develop a hyperparameter-free algorithm for convex composite minimization problems (6) with the optimal theoretical convergence rate $\mathcal{O}(k^{-2})$ while maintaining a reasonable level of computational complexity.
@@
Contributions. Following the same spirit of the adaptive stepsize and linesearch rules, this study proposes a novel stepsize rule without the knowledge of the global smoothness constant for solving convex composite minimization problems. Using a Lyapunov-based argument, we show that the proposed rule enjoys the optimal worst-case complexity bound of $\mathcal{O}(k^{-1})$ and $\mathcal{O}(k^{-2})$ in both cases of nonaccelerated and accelerated algorithms, respectively. These results are developed for a general class of composite convex functions and are optimal in the sense that they match the theoretical lower bounds of the class of first-order algorithms for this class of functions. Our convergence results are summarized in Table 1 where our update direction $d_{k}=-G^{f}_{\lambda h}(x_{k})$ is our gradient mapping (see the definition (10) for more details concerning the gradient mapping). Let us note that the results in Table 1 are reported for the general composite case (6). If the nonsmooth term is $h(x)=0$, then the gradient mapping reduces to $G^{f}_{\lambda h}(x)=\nabla f(x)$. In the rest of the introduction, we briefly discuss our proposed stepsize rule in this special case of a smooth convex function and provide a geometrical illustration in comparison with the existing rules reviewed above.
@@
where the function $\phi_{k}$ is defined in (2) and visualized in Figure 1a. Recall that in the case of non-accelerated methods, the update direction is simply $d_{k}=-\nabla f(x)$. We emphasize that the proposed linesearch above only requires additional zeroth-order information, namely the objective function evaluation, while the first-order information is the gradient of the function which is already computed in the previous step. Figure 1b illustrates a geometric interpretation of the proposed stepsize rule (7), together with the existing literature reviewed in the preceding section. For further geometrical interpretation, particularly in case of accelerated methods, we refer to [bubeck2015geometric].
@@
Notation. Let $\mathcal{X}$ be the finite-dimensional real vector space with the standard inner product $\langle\cdot,\cdot\rangle$ and $\ell_{p}$-norm $\|\cdot\|_{p}$ (by $\|\cdot\|$, we mean the Euclidean standard 2-norm). If $f$ is differentiable, then $\nabla f(x)$ represents the gradient of $f$ at $x$. The function $f$ is $L$-smooth, or equivalently the gradient of $f$ is $L$-Lipschitz, if for all $x,y\in\mathbb{R}^{d}$ one of the following inequalities is satisfied:
@@
Furthermore, $f$ is locally smooth if it is smooth over any compact set of its domain, see [Nesterov2018, Chap. 2] for more details. The prox operator of a convex function $h\colon\mathcal{X}\to\mathbb{R}$ is defined as
@@ 2.1. Preliminaries
We first proceed with some assumptions and lemmas that will be used throughout the paper. We note that the classical Cauchy-Schwartz and convexity inequalities are the two useful tools in our analysis and most of the results build on the seminal works by Nesterov [Nesterov1983] and Polyak [Polyak1969]. The following assumptions hold throughout this study.
@@
[…]
@@
[…]
@@ 2.2. Non-accelerated composite minimization
[…]
Then, we have the uniform stepsize lower bound $\lambda_{k}\geq{1}/{3L}$ where $L$ is the smoothness constant of $f$. Moreover, we have the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D}{k}$ where $D$ is a constant that only depends on the initial condition $x_{0}$, the optimal solution $x^{*}$, and the smoothness constant $L$.
@@
First, to show that $\lambda_{k}$ is bounded from below by 1/3L{1}/{3L}, we just need to show that the condition in the stepsize rule always is satisfied by $\lambda_{k}=1/{3L}$. Considering the smoothness of $f$, we can write the inequality (8b) for the pair $(x,x^{+})$ where $x^{+}=x-\lambda G^{f}_{\lambda h}(x)$, which arrives at the two inequalities
@@
Due to the smoothness of $\phi(\cdot)$, the stepsize rule in Alg. 1 is satisfied whenever
$\big\langle G^{f}_{\lambda h}(x),\,\nabla f(x^{+})-\nabla f(x)+\tfrac{1-L\lambda}{2}\,G^{f}_{\lambda h}(x)\big\rangle\geq 0,$
which coincides with the inequality condition in (14) when $\lambda=1/3L$, showing that the stepsize rule in Alg. 1 is always satisfied by choosing $\lambda=1/3L$. Next, we prove the convergence of Alg. 1. By putting $z=x_{k}$ in the inequality (iii) we have
$F(x_{k+1})-F(x_{k})\leq-\dfrac{\lambda_{k}}{2}\|G^{f}_{\lambda_{k}h}(x_{k})\|^{2}$ (15)
Telescoping the above inequalities from $k=0$ to $k=\infty$ and using the fact that $\lambda_{k}\geq 1/3L$ gives us
$F(x^{\infty})-F(x_{0})\leq-\dfrac{1}{6L}\sum_{i=0}^{\infty}\|G^{f}_{\lambda_{i}h}(x_{i})\|^{2}$
@@
The above approximate solution is guaranteed to satisfy $\lambda_{k}\geq\dfrac{C}{3L}$.
@@ 2.3. Non-accelerated smooth minimization
Let the function $F$ in (6) be a smooth (i.e., $h(x)=0$). Then, the stepsize rule in Alg. 1 reduces to (7) while ensuring Theorem 2.3 results of the uniform stepsize lower bound $\lambda_{k}\geq{1}/{3L}$ and the sub-optimality bound $F(x_{k})-F(x^{*})\leq\dfrac{D}{k}$.
@@
We close this subsection by highlighting a common feature shared between the Polyak stepsize (4), Armijo stepsize rule (3a), and our proposed stepsize rule (7). These three stepsizes, denoted by $\lambda$, respect the following inequalities
$\dfrac{1}{3L}<\dfrac{1}{2L}\leq\lambda\leq c\,\dfrac{f(y_{1})-f(y_{2})}{\|\nabla f(x_{k})\|^{2}},$ (18a)
@@
The common lower bound $1/(2L)$ in (18a) is well known for Polyak- and Armijo-type rules. Our proposed rule in Theorem 2.3 instead achieves a lower bound of $1/(3L)$, which is smaller than the bound of the existing linesearch rules. While larger stepsizes are commonly expected to yield faster convergence, our numerical experiments across multiple problem classes show that the proposed linesearch consistently outperforms existing methods. A possible explanation for this is that the theoretical bounds are typically conservative: in realistic problems, the local geometry often permits stepsizes much larger than those implied by the global smooth constant. Our linesearch procedure exploits this favorable curvature by accepting substantially larger steps in most iterations, leading to faster objective decrease and fewer gradient/proximal evaluations overall.
Notice that the optimal choice of the stepsizes is the upper bound of the admissible interval (18a). It is also important to note that in the case of Armijo (3a) and our proposed choice (7), this upper bound also depends on $\lambda$ (see (18b)), making the stepsize rule implicit. Finally, we wish to draw attention to the choice of the upper bound between Armijo and our proposed approach in (18a): One can inspect that the proposed rule (7) looks one step ahead of Armijo in the direction of $\nabla f$. This may explain why the proposed rule (7) performs numerically better, as it can approximate the curvature of the function $f$ by “looking one step further” than Armijo. Finally, it is also worth noting that the recent work [Malitsky2020] offers an adaptive stepsize rule with the lower bound $1/2L$ and an explicit upper bound (i.e., independent of $\lambda$) that depends on the gradient of two successive iterations (see (5)).